For each of the following independent situations determine the amount of the standard deduction that the taxpayers should claim on their 2018 income tax returns.

Table for the standard deduction

Filing Status 2018 Standard Deduction
Single $ 12,000
Married, filing jointly 24,000
Married, filing separately 12,000
Head of household 18,000
Qualifying widow(er) 24,000
1. Adam is 45 years old, in good health, and single. $
2. Bill and Betty are married and file a joint return. Bill is 66 years old, and Betty is 60. $
3. Charlie is 70, single, and blind. $
4. Debbie qualifies for head of household filing status, is 35 years old, and is in good health. $
5. Elizabeth is 9 years old, and her only income is $3,600 of interest on a savings account. She is claimed as a dependent on her parents' tax return. $
6. Frank and Frieda are married with two dependent children. They file a joint return, are in good health, and both of them are under 65 years of age. $
Answer #1

1) $12,000

2)24,000 +1,300 =25,300

3)12,000+1,600+,1600= 15,200

4)18,000

5)3,600+350=3,950

6)24,000